The sum of two consecutive even integers is equal to 30. What is the smaller integer?
Answer by mananth(16946)

let the numbers be x & x+2
x+x+2=30
2x+2=30
2x=30-2
2x=28
x=28/2
x=14
..
The numbers are 14 & 16